OpenStudy (anonymous):

(ANYONE good at probabilities?)The train station runs trains every 30 minutes. Find the probability that a passenger who arrives at a random time will not have to wait more than 20 minutes. A.1/3 B.1/2 C.2/3 D.25/30

OpenStudy (anonymous):

A or B. * oops.

OpenStudy (anonymous):

hello??

OpenStudy (jack1):

i wanna say c... *train arrives 1...>20mins 2...>20mins 3...>20mins 4...>20mins 5...>20mins 6...>20mins 7...>20mins 8...>20mins 9...>20mins up until exactly 10 ...>20mins 11...<20mins 12...<20mins 13...<20mins 14...<20mins 15...<20mins 16...<20mins 17...<20mins 18...<20mins 19...<20mins 20...<20mins 21...<20mins 22...<20mins 23...<20mins 24...<20mins 25...<20mins 26...<20mins 27...<20mins 28...<20mins 29...<20mins *train arrives sum is 10/30 @...>20mins 20/30 @...<20mins

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

Thanks for typing all of that out. x)

OpenStudy (jack1):

yeah P("will not have to wait more than 20 minutes") = 1- P(will have to wait more than 20 minutes) = 1 - 1/3 = 2/3

OpenStudy (jack1):

s ók

OpenStudy (jack1):

oooo shiny